solve 6x-4(x+1)= 2(x-3)+2
Answer by stanbon(75887)

solve 6x-4(x+1)= 2(x-3)+2
6x-4x-4 = 2x-6+2
2x-4 = 2x-4
------------
That is true for all values of "x".
The original problem is an identity.
=====================================
Cheers,
Stan H.